# FlashSinkhorn: IO-Aware Entropic Optimal Transport

### v1

@@ G.2 Loop order
5:  Load $K_{J}$ to on-chip SRAM
@@
8:   Load $Q_{I}$, $\hat{\mathbf{f}}_{I}$ and $\bm{\gamma}_{I}$ to on-chip SRAM
9:   $S\leftarrow(K_{J}Q_{I}^{\top}+\hat{\mathbf{f}}_{I}+\bm{\gamma}_{I})/\varepsilon$ // scores + bias
@@
5:  Load $K_{J}$, $\hat{\mathbf{g}}_{J}$, $\mathbf{b}_{J}$ to on-chip SRAM
6:  $\mathbf{m}_{J}\leftarrow$-$\infty$, $\mathbf{s}_{J}\leftarrow 0$, $O_{J}\leftarrow 0$
@@
8:   Load $Q_{I}$, $\hat{\mathbf{f}}_{I},\bm{\gamma}_{I}$, $V_{I}$ to on-chip SRAM
9:   $S\leftarrow(K_{J}Q_{I}^{\top}+\hat{\mathbf{f}}_{I}+\bm{\gamma}_{I})/\varepsilon$ // score tile + bias
@@
13:   $O_{J}\leftarrow e^{\mathbf{m}_{J}-\mathbf{m}_{\mathrm{new}}}\odot O_{J}+e^{S-\mathbf{m}_{\mathrm{new}}}\,V_{I}$
14:   $\mathbf{m}_{J}\leftarrow\mathbf{m}_{\mathrm{new}}$
15:  end for
16:  $\hat{\mathbf{g}}^{+}_{J}\leftarrow-\varepsilon\cdot\big(\mathbf{m}_{J}+\log\mathbf{s}_{J}\big)$ // one-step $\hat{\mathbf{g}}-$update
17:  $\mathbf{c}_{J}\leftarrow\mathbf{b}_{J}\odot\exp\!\left((\hat{\mathbf{g}}_{J}-\hat{\mathbf{g}}^{+}_{J})/\varepsilon\right)$ // $\mathbf{c}=(P^{\top}\mathbf{1})_{J}$
18:  $\mathrm{out}_{J}\leftarrow\mathrm{diag}(\mathbf{c}_{J})\,\mathrm{diag}(\mathbf{s}_{J})^{-1}\,O_{J}$
19:  Write $\mathrm{out}_{J}$ to HBM
20:end for
Algorithm 5 Hadamard-Weighted Transport from Potentials, $(P\odot W)\ V$ with $W_{kj}=A_{k}B_{j}^{\top}$
@@
5:  Load $Q_{I}$, $A_{I}$, $\hat{\mathbf{f}}_{I}$, $\mathbf{a}_{I}$ to on-chip SRAM
6:  $\mathbf{m}_{I}\leftarrow$-$\infty$, $\mathbf{s}_{I}\leftarrow 0$, $O_{I}\leftarrow 0$ // running max/sumexp, output accumulator
@@
8:   Load $K_{J}$, $B_{J}$, $\hat{\mathbf{g}}_{J},\bm{\delta}_{J}$, $V_{J}$ to on-chip SRAM
9:   $S\leftarrow(Q_{I}K_{J}^{\top}+\hat{\mathbf{g}}_{J}+\bm{\delta}_{J})/\varepsilon$ // score tile + bias
10:   $W\leftarrow A_{I}B_{J}^{\top}$ // Hadamard weights: $W_{kj}=A_{k}B_{j}^{\top}$
11:   $\tilde{\mathbf{m}}\leftarrow\mathrm{rowmax}(S)$
@@
13:   $\mathbf{s}_{I}\leftarrow e^{\mathbf{m}_{I}-\mathbf{m}_{\mathrm{new}}}\odot\mathbf{s}_{I}+\mathrm{rowsum}(e^{S-\mathbf{m}_{\mathrm{new}}})$
14:   $O_{I}\leftarrow e^{\mathbf{m}_{I}-\mathbf{m}_{\mathrm{new}}}\odot O_{I}+(e^{S-\mathbf{m}_{\mathrm{new}}}\odot W)\,V_{J}$
15:   $\mathbf{m}_{I}\leftarrow\mathbf{m}_{\mathrm{new}}$
16:  end for
17:  $\hat{\mathbf{f}}^{+}_{I}\leftarrow-\varepsilon\cdot\big(\mathbf{m}_{I}+\log\mathbf{s}_{I}\big)$ // one-step $\hat{\mathbf{f}}-$update
18:  $\mathbf{r}_{I}\leftarrow\mathbf{a}_{I}\odot\exp\!\left((\hat{\mathbf{f}}_{I}-\hat{\mathbf{f}}^{+}_{I})/\varepsilon\right)$ // $\mathbf{r}=(P\mathbf{1})_{I}$
19:  $\mathrm{out}_{I}\leftarrow\mathrm{diag}(\mathbf{r}_{I})\,\mathrm{diag}(\mathbf{s}_{I})^{-1}\,O_{I}$
20:  Write $\mathrm{out}_{I}$ to HBM
21:end for

### v2

@@ G.2 Loop order
5:  Load $Y_{J}$ to on-chip SRAM
@@
8:   Load $X_{I}$, $\hat{\mathbf{f}}_{I}$ and $\bm{\gamma}_{I}$ to on-chip SRAM
9:   $S\leftarrow(2Y_{J}X_{I}^{\top}+\hat{\mathbf{f}}_{I}+\bm{\gamma}_{I})/\varepsilon$ // score tile + source bias
@@
5:  Load $Y_{J}$, $\hat{\mathbf{g}}_{J}$, $\mathbf{b}_{J}$ to on-chip SRAM
6:  $\mathbf{m}_{J}\leftarrow$-$\infty$, $O_{J}\leftarrow 0$
@@
8:   Load $X_{I}$, $\hat{\mathbf{f}}_{I}$, $\bm{\gamma}_{I}$, $V_{I}$ to on-chip SRAM
9:   $S\leftarrow(2\,Y_{J}X_{I}^{\top}+\hat{\mathbf{f}}_{I}+\bm{\gamma}_{I})/\varepsilon$ // score tile + source bias
@@
12:   $O_{J}\leftarrow e^{\mathbf{m}_{J}-\mathbf{m}_{\mathrm{new}}}\odot O_{J}+e^{S-\mathbf{m}_{\mathrm{new}}}\,V_{I}$ // online weighted sum
13:   $\mathbf{m}_{J}\leftarrow\mathbf{m}_{\mathrm{new}}$
14:  end for
15:  $\mathrm{out}_{J}\leftarrow\mathbf{b}_{J}\odot\exp\!\left(\hat{\mathbf{g}}_{J}/\varepsilon+\mathbf{m}_{J}\right)\odot O_{J}$ // target marginal correction
16:  Write $\mathrm{out}_{J}$ to HBM
17:end for
Algorithm 5 Hadamard-Weighted Transport from Potentials, $(P\odot W)\,V$ with $W_{ij}=A_{i}B_{j}^{\top}$
@@
5:  Load $X_{I}$, $A_{I}$, $\hat{\mathbf{f}}_{I}$, $\mathbf{a}_{I}$ to on-chip SRAM
6:  $\mathbf{m}_{I}\leftarrow$-$\infty$, $O_{I}\leftarrow 0$ // running max, Hadamard-weighted accumulator
@@
8:   Load $Y_{J}$, $B_{J}$, $\hat{\mathbf{g}}_{J}$, $\bm{\delta}_{J}$, $V_{J}$ to on-chip SRAM
9:   $S\leftarrow(2\,X_{I}Y_{J}^{\top}+\hat{\mathbf{g}}_{J}+\bm{\delta}_{J})/\varepsilon$ // score tile + target bias
10:   $W\leftarrow A_{I}B_{J}^{\top}$ // Hadamard weight tile
11:   $\tilde{\mathbf{m}}\leftarrow\mathrm{rowmax}(S)$ // tile max
@@
13:   $O_{I}\leftarrow e^{\mathbf{m}_{I}-\mathbf{m}_{\mathrm{new}}}\odot O_{I}+(e^{S-\mathbf{m}_{\mathrm{new}}}\odot W)\,V_{J}$ // online weighted sum
14:   $\mathbf{m}_{I}\leftarrow\mathbf{m}_{\mathrm{new}}$
15:  end for
16:  $\mathrm{out}_{I}\leftarrow\mathbf{a}_{I}\odot\exp\!\left(\hat{\mathbf{f}}_{I}/\varepsilon+\mathbf{m}_{I}\right)\odot O_{I}$ // source marginal correction
17:  Write $\mathrm{out}_{I}$ to HBM
18:end for